# UltraVSR: Achieving Ultra-Realistic Video Super-Resolution with Efficient One-Step Diffusion Space

### v1

@@ 4.3. Ablation Study
We conduct comprehensive quantitative and qualitative experiments to evaluate the contributions of the key components in UltraVSR. The ablation results are presented in Tab. 3, with exemplar visual comparisons shown in Fig. 7. We begin with Model A, which serves as our baseline, constructed by: (1) fixing the diffusion time step to 999, following the one-step generation paradigm established in prior works (Yin et al., 2024; Nguyen and Tran, 2024; Frans et al., 2024; Xie et al., 2024); (2) enabling only the trainable LoRA layers; and (3) replacing our proposed Spatio-temporal Joint Distillation (SJD) with the conventional Variational Score Distillation (VSD) (Yin et al., 2024; Wang et al., 2023b). Due to its inability to dynamically perceive degradation levels, Model A exhibits limited reconstruction quality compared to Model B (with DRS). Next, we progressively enhance Model B by introducing trainable RTS-convolution and RTS-attention units, resulting in Model C and Model D, respectively. Thanks to the strong temporal modeling capability of RTS, inter-frame consistency improves steadily across these models. We further evaluate the effectiveness of SJD and Temporally Asynchronous Inference (TAI) by comparing Model D, Model E (where a fixed local sequence of length 10 is used), and our model. It can be seen that SJD significantly enhances temporal coherence by introducing inter-frame difference supervision, leading to improved motion continuity across frames compared to VSD (Yin et al., 2024; Wang et al., 2023b). Moreover, TAI further boosts performance by enabling long-range temporal modeling. These results collectively demonstrate the necessity and effectiveness of each proposed component in UltraVSR.
5. Discussion
While UltraVSR introduces an efficient one-step VSR framework that greatly reduces the computational cost, it is still challenging to deploy it on multimedia applications with high real-time requirements. We will explore hardware-aware acceleration on different deployment platforms in future work.
@@ 6. Conclusion
We present UltraVSR, the first work to address real-world VSR through an efficient one-step diffusion framework. To enable this, we propose Degradation-aware Reconstruction Scheduling (DRS), which transforms the conventional multi-step denoising into a one-step LR-to-HR reconstruction. To maintain temporal consistency, we develop the Recurrent Temporal Shift (RTS) module for efficient feature propagation, fusion, and alignment. We further introduce Spatio-temporal Joint Distillation (SJD) for realistic and coherent outputs, and Temporally Asynchronous Inference (TAI) for efficient long-range temporal modeling. Extensive experiments show that UltraVSR delivers favorable performance with significantly accelerated inference speed.

### v2

@@ 4.3. Ablation Study
We conduct comprehensive experiments to evaluate the contributions of the key components in UltraVSR. The results are presented in Tab. 3 and Fig. 7. We begin with Model A, which serves as our baseline, constructed by: (1) fixing the diffusion time step to 999, following the one-step generation paradigm established in prior works (Yin et al., 2024; Nguyen and Tran, 2024; Frans et al., 2024; Xie et al., 2024); (2) enabling only the trainable LoRA layers; and (3) replacing our proposed Spatio-temporal Joint Distillation (SJD) with the conventional Variational Score Distillation (VSD) (Yin et al., 2024; Wang et al., 2023b). Due to its inability to dynamically perceive degradation levels, Model A exhibits limited reconstruction quality compared to Model B (with DRS). We incrementally enhance Model B with trainable RTS-convolution and RTS-attention units to form Models C and D, respectively, leading to steadily improved inter-frame consistency. We further evaluate the effectiveness of SJD and TAI by comparing Model D, Model E (with a fixed local sequence of length 10), and our model. It can be seen that SJD significantly enhances temporal coherence by introducing inter-frame difference supervision, leading to improved motion continuity across frames compared to VSD (Yin et al., 2024; Wang et al., 2023b). TAI further boosts performance by enabling long-range temporal modeling.
@@ 5. Conclusion
We present UltraVSR, the first real-world VSR framework based on an efficient one-step diffusion space. It introduces Degradation-aware Reconstruction Scheduling (DRS) to transform multi-step denoising with one-step LR-to-HR reconstruction. To maintain temporal consistency, we develop the Recurrent Temporal Shift (RTS) module and Spatio-temporal Joint Distillation (SJD) for realistic and coherent outputs, along with Temporally Asynchronous Inference (TAI) for efficient long-range temporal modeling. Extensive experiments show that UltraVSR delivers favorable performance with significantly accelerated inference speed.
6. Acknowledgments
This work was supported in part by the National Major Science and Technology Projects of China under Grant 2009XJTU0016, in part by the National Natural Science Foundation of China under Grant U22B2049.